### v1

# MS-Temba : Multi-Scale Temporal Mamba for Efficient Temporal Action Detection

Abstract
Action detection in real-world scenarios is particularly challenging due to densely distributed actions in hour-long untrimmed videos. It requires modeling both short- and long-term temporal relationships while handling significant intra-class temporal variations. Previous state-of-the-art (SOTA) Transformer-based architectures, though effective, are impractical for real-world deployment due to their high parameter count, GPU memory usage, and limited throughput, making them unsuitable for very long videos. In this work, we innovatively adapt the Mamba architecture for action detection and propose Multi-scale Temporal Mamba (MS-Temba), comprising two key components: Temporal Mamba (Temba) Blocks and the Temporal Mamba Fuser. Temba Blocks include the Temporal Local Module (TLM) for short-range temporal modeling and the Dilated Temporal SSM (DTS) for long-range dependencies. By introducing dilations, a novel concept for Mamba, TLM and DTS capture local and global features at multiple scales. The Temba Fuser aggregates these scale-specific features using Mamba to learn comprehensive multi-scale representations of untrimmed videos. MS-Temba is validated on three public datasets, outperforming SOTA methods on long videos and matching prior methods on short videos while using only one-eighth of the parameters.  [^1] The code is available at https://github.com/thearkaprava/MS-Temba
Figure 1: Temporal Action Detection using Mamba: On the left, a visual sample demonstrates that our model, MS-Temba, accurately localizes detected action labels in comparison to MS-TCT for temporal action detection in untrimmed videos. On the right, a plot of GPU memory usage versus video duration highlights the efficiency of MS-Temba in processing long untrimmed videos.
“Temba, his arms wide.”
Star Trek: The Next Generation episode,“Darmo”
1 Introduction
Learning representations in long, untrimmed videos is challenging due to complex, long-term dependencies that advanced neural networks struggle to optimize end-to-end. One such task, action detection, involves precisely localizing temporal boundaries of actions in untrimmed videos, with applications in areas like patient monitoring and video surveillance. Real-world deployment of temporal action detection models requires efficiency in capturing actions and their boundaries, yet real-world videos differ significantly from sports datasets [22, 2], where a few, non-overlapping actions with clear motion patterns are present. Real-world videos often contain densely distributed actions over time with subtle motion patterns, further complicated by large intra-class variance (e.g., the same action occurring with varying durations), the need for long-term relation modeling, and simultaneous actions. This paper focuses on developing an efficient model for action detection in hour-long videos, addressing these real-world challenges.
To learn temporal representations in untrimmed videos, early methods employed proposal-based approaches [43], which, while effective for videos with sparse action distributions, are resource-intensive due to the need to generate numerous proposals. Consequently, temporal convolutional methods [14, 31, 32] became popular for action detection, although their local receptive fields limit the capture of long-term dependencies. With the success of self-attention mechanisms, transformer-based methods [39, 8] were introduced to learn global representations across time frames. Despite their effectiveness, these transformer models impose high computational costs in terms of model parameters, memory usage, and throughput, making them challenging to deploy on edge devices for real-world applications.
In recent research, Mamba [17, 26, 50] is gaining attention for its efficiency in handling long sequences with linear complexity, making it well-suited for action detection in untrimmed videos. While Mamba has seen exploration in the vision domain, its application to action detection in untrimmed videos remains underexplored, motivating us to develop a Mamba-based architecture tailored for this purpose. To address this, we propose Multi-scale Temporal Mamba (MS-Temba), an architecture designed for efficient action detection (see Figure 1). MS-Temba incorporates two key components: 1) Temporal Mamba Blocks (Temba blocks) and 2) the Temporal Mamba Fuser (Temba fuser). The Temba Blocks are designed to learn temporal representations at multiple scales, achieved through two modules: the Temporal Local Module (TLM) and the Dilated Temporal State Space Module (DTS). TLM captures local temporal representations, while DTS learns long-range relationships using state space models enhanced with dilations. Although dilated convolutions are widely used, this work is the first to incorporate dilations into SSMs, allowing DTS to skip immediate neighbors and capture relationships across different temporal scales. This approach enhances Temba Blocks’ ability to model long-term temporal relations effectively. Additionally, the Temba Fuser aggregates scale-specific temporal information from each Temba Block, learning a multi-scale representation. This fusion, performed via Mamba, is computationally lighter than dense layers. We assess MS-Temba on three publicly available datasets with densely labeled actions, finding that it performs on-par with previous state-of-the-art method on short-duration videos and surpasses it on long-duration videos, using only one-eighth the parameters. Our study reveals that MS-Temba outperforms transformer-based models when both use a foundational visual backbone, highlighting Mamba’s potential not only for long sequences but also for feature inputs from foundation models [33].
To summarize the main contributions:
  1. 1.
    We introduce a novel Mamba-based architecture Multi-scale Temporal Mamba (MS-Temba), tailored for densely labeled Temporal Action Detection
  2. 2.
    MS-Temba explores different temporal scales through our novel components in Temba block, especially dilated Temporal SSMs, and aggregates them to learn multiscale representation through the Temporal Mamba Fuser
  3. 3.
    This carefully crafted design of MS-Temba  achieves performance on par with Transformer-based architectures on short videos and outperforms the previous state-of-the-art on long-duration videos with $88\%$ less parameters.
Figure 1 illustrates a visual sample of action detection using MS-Temba compared to previous state-of-the-art methods. MS-Temba efficiently processes hour-long videos with low GPU memory usage and fewer parameters, maintaining high accuracy. This motivates the suitability of the Mamba architecture for temporal action detection in untrimmed videos.
2 Related Work
Understanding actions in long temporal sequences has become a key task in computer vision [16, 27, 44, 11, 10, 48]. Early action detection approaches focused on videos with sparse action labels [22, 48, 2], primarily using proposal-based methods [43, 49] inspired by object detection to generate action proposals. However, temporal action detection with densely labeled action distribution offers greater challenges due to concurrent actions, where proposal-based methods struggle with the combinatorial explosion in proposal generation.
This led to the adoption of temporal convolution-based models [31, 14] for processing long videos. Convolutional methods like PDAN [7] and TGM [32] use specialized kernels designed to model long-term dependencies and capture composite actions. However, the shared kernels in temporal convolutions are limited to neighboring frames, and even with dilations, struggle to capture global relationships between frames. This limitation is significant for densely labeled Temporal Action Detection, where long-term actions are common. In response, transformer-based architectures [41, 12, 40, 29, 13], which have proven effective in computer vision, have been adapted for action detection [39, 8, 34]. These architectures can model both short- and long-term dependencies in videos. For example, MLAD [39] employs a Transformer encoder to capture cross-class features and temporal relationships, while MS-TCT [8] uses a temporal hierarchy to model local relationships through temporal convolutions and global interactions through attention. These transformer based approaches enhances multi-scale representation learning, but the effectiveness comes at a high computational cost. The linear complexity of state space models (Mamba [17]) makes it an ideal test bed for processing long untrimmed videos.
In recent times, Structured State Space Model (S4) [18] has gained popularity in effectively processing long sequence lengths in linear complexity. Mamba [17] addresses some challenges of S4 (selective copying and induction heads) and introduces data-dependent SSM layer and a hardware aware efficient algorithm. Consequently, Mamba has been has been adapted for various computer vision downstream tasks using static images such as image classification, object detection, semantic segmentation, image restoration in [50, 51, 19, 20, 42, 28, 1]. Naturally, these state space models have been adapted for videos in a wide range of applications such as action classification [26, 52], video object segmentation [45], video retrieval [38] and physiological measurement from face videos [30]. However, these methods are trained on videos which span only few seconds and have a global class per video. It is non trivial to adapt these methods for action detection since action detectors requires two stage training, first feature extraction from untrimmed videos using a visual backbone and then temporal learning of these features for dense action-label prediction. For long-range temporal modeling, TranS4mer [21] has shown effectiveness in video classification and movie scene detection tasks by identifying scene changes within videos. Closest to our approach, Video Mamba Suite [4] performs action localization on datasets like HACS [48] and GTEA [15]. These Mamba-based methods rely on frame sampling and compressing all information within their architectures limiting them to process videos around 3 minutes. However, in this work, we target untrimmed videos exceeding 40 minutes, featuring dense labels and co-occurring actions. Linear-complexity state space models have not yet been applied to action detection in such untrimmed videos. To the best of our knowledge, MS-Temba is the first to leverage Mamba for the task of temporal action detection in densely labeled untrimmed videos.
3 Preliminaries
State space models map continuous input sequences $x(t)\in\mathbb{R}$ to continuous outputs $y(t)\in\mathbb{R}$ using a hidden state $h(t)\in\mathbb{R}^{n}$. For the above mapping and updating of the hidden state, SSMs use a transition state matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ projection matrices $\mathbf{B}\in\mathbb{R}^{n\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times n}$. The update equations are given by
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{A}x(t)$ (1)
$y(t)$ $=\mathbf{C}h(t)$
Previous works  [18, 17, 50], use discrete versions of the above system using a discretization parameter $\mathbf{\Delta}$ known as the step size. Following is an example of a discretization transformation using $\mathbf{\Delta}$, known as zero-order hold (ZOH):
$\overline{\mathbf{A}}$ $=\exp(\mathbf{\Delta}\mathbf{A})$ (2)
$\overline{\mathbf{B}}$ $=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp(\mathbf{\Delta}\mathbf{A})-\mathbf{I})\mathbf{\Delta}\mathbf{B}$
This discretization of $\mathbf{A}$, $\mathbf{B}$ to $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ leads to the following updated state-space equations:
$h_{t}$ $=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t}$ (3)
$y_{t}$ $=\mathbf{C}h_{t}$
Standard SSMs are typically designed to process 1D inputs in a linear sequence. However, [50, 26] introduce bidirectional scanning of tokens (illustrated in Figure 3), which are more effective for vision tasks. In this approach, the SSM operates on the input sequence $v$ and it’s reversed counterpart $v^{\prime}$, fusing the projected outputs for robust bidirectional sequence modeling. For brevity, we call it BiMamba for the remainder of the paper. This enhancement enables more effective representation learning, especially in domains requiring complex temporal dependencies. In this work, we extend the Mamba based architectures for understanding long untrimmed videos.
4 Multi-scale Temporal Mamba
Figure 2: Overall Architecture of Multiscale Temporal Mamba (MS-Temba) for action detection. MS-Temba is composed of a Visual Backbone, a Temporal Encoder consisting of Temporal Mamba (Temba) Blocks, a Temporal Mamba Fuser and a Classification Head.
In Temporal Action Detection, we aim to classify multiple activities at each time-step of an untrimmed video. Formally, given a video of $T$ temporal segments, we associate an action label $y^{c}_{t}$ for each time-step $t$, where $c$ denotes the action classes, with $c\in\{1,...,C\}$.
Here, we introduce Multi-Scale Temporal Mamba (MS-Temba), which is capable of detecting fine-grained as well as composite actions efficiently in long untrimmed videos. As illustrated in Figure 2, MS-Temba  consists of: i) a Visual Backbone to extract discriminative features from untrimmed videos, ii) Temba blocks which captures temporal information at different scales, iii) Temba Fuser to combine the learned representations across different scales, and iv) a Classification Head to predict multiple actions within each temporal segment of the input video. In the following sections, we detail each of these components of MS-Temba.
4.1 Visual Backbone
The input to MS-Temba is an untrimmed video potentially spanning long durations (several minutes). Training an end-to-end model across both spatial and temporal dimensions poses significant computational demands. Consequently, in line with prior work [8, 7, 14, 47], we first segment the untrimmed video into $T$ contiguous, non-overlapping segments, each consisting of 8 or 16 frames. Each segment, known as a visual token, is encoded via a visual backbone, capturing spatial features through either a 3D ConvNet (e.g., I3D [3]) pre-trained on the target dataset, or a larger foundational model such as CLIP [33]. Unlike action classification tasks, sparse frame sampling is unsuitable here, as it risks omitting short-duration actions crucial for detection. The segment-wise representations are then stacked to form an input feature vector $v\in\mathbb{R}^{T\times D}$, where D is the feature dimension and T is the total temporal sequence length. Finally, $v$ is projected to $z_{0}\in\mathbb{R}^{T\times D_{0}}$ to obtain a latent representation of the input, which serves as input tokens to our proposed Mamba-based temporal encoder.
Figure 3: Temba Block. A single Temba block consists of a Temporal Local Module (TLM) and a Dilated Temporal SSM. The TLM is responsible for learning local feature while the DTS learns long-range temporal dependencies. $\Phi$ denotes a non-parameterized operation to facilitate dilation in the SSM.
4.2 Temporal Mamba (Temba) Blocks
For understanding the semantics of short-term and long-term actions in untrimmed videos, it is essential to model temporal relations at different scales. To this end, we introduce Temporal Mamba (Temba) Blocks which consists of two primary modules. i) Temporal Local Module to capture local temporal relations between tokens and ii) Dilated Temporal SSM to capture long-range temporal dependencies within videos.
Temporal Local Module (TLM) forms an essential component of Temba blocks, specifically designed to introduce local interactions between tokens in neighboring temporal region. In order to capture multiscale temporal representations it is important to process the tokens locally at different neighborhoods. Therefore, TLM uses a dilated temporal convolution layer with a kernel size of $k=3$, stride of $1$ and dilation of $\delta$. We gradually increase $\delta$ in subsequent blocks. This layer incrementally captures temporally spaced features, allowing the model to focus on local neighborhoods in early blocks, where the dilation is minimal, and progressively incorporates more distant neighborhoods in later blocks as the dilation increases. Formally, each module can be described as follows:
$z_{b}=W_{b}z_{b-1}+\beta_{b},\quad z_{b}\in\mathbb{R}^{B\times T\times D_{b}}$ (4)
where $W_{b}$ is the 1D convolutional kernel of the $b^{th}$ Temba Block and $\beta_{b}$ is the bias, $D_{b}$ is the feature dimension computed as $D_{b}=\gamma D_{b-1}$. Thus, the TLM progressively increases the feature dimension by a factor of $\gamma$ in each successive block, enabling the model to capture complex patterns effectively as its receptive field increases. The receptive field for a kernel of size 3 and dilation $\delta$ is given by,
$ReceptiveField(\delta)=2^{\delta}+1$ (5)
Functionalities of the TLM can also be achieved using MLPs or more complex networks for local interaction. Consequently, the TLM introduces a strong inductive bias by mapping token representations across multiple temporal scales, subsequently enabling the temporal SSM to capture global token interactions effectively.
Figure 4: Scanning Strategy. An illustration of scanning strategy in Dilated Temporal SSM (DTS). Note this diagram shows only the forward scanning in the second Temba Block. MS-Temba uses bidirectional scanning.
Dilated Temporal SSM. (DTS) While the TLM captures interactions in a local neighborhood, it is not sufficient for learning long-range temporal relationships among actions which is crucial for Temporal Action Detection, as discussed in Section 1. These global interactions are effective in Transformer-based architectures, but they come at a high computational cost and are infeasible for hour-long videos. However, SSMs, albeit efficient in processing long sequences, scan the input tokens sequentially and update the state-space only with neighboring token features. We aim to introduce the concept of dilation in SSMs, to effectively capture the long-range relationships. To this end, we introduce a non-parameterized invertible operation $\mathit{\Phi}_{\eta}:\mathbb{R}^{B\times T\times D}\rightarrow\mathbb{R}^{\eta B\times T/\eta\times D}$, which is responsible for extracting distinct tokens at different temporal resolutions controlled by the SSM dilation rate $\eta$. For example, for $\eta=2$, the temporal sequence $T=[t_{1},t_{2},t_{3},t_{4},..,t_{2N}]$ is divided into two subsets $T_{1}$ and $T_{2}$, consisting of non-overlapping tokens $T_{1}=[t_{1},t_{3},t_{5},...,t_{2N-1}]$ and $T_{2}=[t_{2},t_{4},t_{6},...,t_{2N}]$, where $T_{1}\cap T_{2}=\phi$. These are scanned separately by a bidirectional SSM (BiMamba), illustrated in Figure 4, and the learned representations are merged together in the respective initial positions (as in $T$) to achieve a global interaction among the tokens in the SSM. This approach, dubbed Dilated Temporal SSM, is formally described as follows:
$z_{b}$ $=\mathit{\Phi}_{\eta}(z_{b}),\quad z_{b}\in\mathbb{R}^{\eta B\times T/\eta\times D}$ (6)
$z_{b}$ $=BiMamba(z_{b}),\quad z_{b}\in\mathbb{R}^{\eta B\times T/\eta\times D}$ (7)
$z_{b}$ $=\mathit{\Phi}^{-1}_{\eta}(z_{b}),\quad z_{b}\in\mathbb{R}^{B\times T\times D}$ (8)
The combination of TLM with dilated token scanning enables the Temba blocks to refine state representations across layers at various temporal scales, defined by the two dilation parameters, $\delta$ and $\eta$.
Figure 5: Temporal Mamba Fuser (TMF) is composed of projection layers to aggregate features learnt at different Temba Blocks. After these features are fused, BiMamba is responsible for interaction between the fused features.
4.3 Temporal Mamba Fuser
The Temoral Mamba Fuser (Temba Fuser), illustrated in Figure 5, plays a crucial role in aggregating temporal features learned across different scales through Temba blocks, enabling the learning of multi-scale temporal representations. It first applies a linear projection to the output of each block standardizing the feature dimensionality to match that of the final block ($E$). The resulting features are fused to form a unified temporal representation. Finally, we leverage a BiMamba to facilitate interactions across the fused representation from various blocks, each capturing distinct temporal patterns, as follows:
$z_{f}$ $=\sum_{b}W_{f}z_{b}+\beta_{f},\quad z_{f}\in\mathbb{R}^{B\times T\times E}$ (9)
$y^{\prime}$ $=BiMamba(z_{f}),\quad y^{\prime}\in\mathbb{R}^{B\times T\times E}$ (10)
Here, $W_{f}$ and $\beta_{f}$ denote the weight matrix and bias of the linear projector.
4.4 Classification Head
The Classification Head receives the fused features from the Temba Fuser post interaction in the BiMamba. This component is designed as a linear layer that projects the high-dimensional temporal features into the space of multi-label action classes.
$\tilde{y}=W_{h}y^{\prime}+B_{h},\quad\tilde{y}s\in\mathbb{R}^{B\times T\times C}$ (11)
Here, $y^{\prime}\in\mathbb{R}^{B\times T\times E}$ is the representation obtained from Temba Fuser, $W_{h}$ and $B_{h}$ are the parameters of the classification head projector and C is the number of action classes. To train the model effectively for multi-label action detection, following [8, 39], we optimize MS-Temba using a Binary Cross-Entropy (BCE) loss, given by
$\mathcal{L}=-\frac{1}{B}\sum_{i=1}^{B}\left[y_{i}\log(\tilde{y}_{i})+(1-y_{i})\log(1-\tilde{y}_{i})\right]$ (12)
where B is the number of samples in a mini-batch and $y_{i}$ is the ground truth label.
5 Experiments
Datasets. We evaluate MS-Temba, on three multi-label action detection datasets: Toyota Smarthome Untrimmed [5], MultiTHUMOS [46], and Charades [35]. All three datasets contain densely labeled activities, making the task of action detection challenging. Toyota Smarthome Untrimmed (TSU) [5] is an untrimmed video dataset capturing indoor activities related to daily living. It includes 51 action classes, with up to 5 co-occurring actions in a single frame, as well as composite actions. The average duration of the videos in TSU is around 21 minutes, considerably higher than the other datasets making it challenging for action detection. MultiTHUMOS [46] is another densely annotated dataset containing frame-level annotations of 400 videos from the THUMOS’14 [22] dataset. It consists of 65 action classes with an average of 10.5 action instances per video and 1.5 labels per frame. Charades [35] is another large-scale dataset consisting of 9848 untrimmed videos spanning 157 actions. It contains 66500 annotations of daily activities with an average of 6.8 action instances per video. We evaluate our methods on these datasets using the ”action localization” protocol described in [5][36], and [46], respectively, to detect frame-level actions and compare them with the temporal annotations.
Implementation Details. In MS-Temba, we use I3D [3] and CLIP-L/14 [33] as our visual backbones. Their feature dimensions $D$ are 1024 and 768 respectively. The input features are projected to $D_{0}=256$ before it is input to the first Temba block. MS-Temba consists of 3 Temba blocks with a feature expansion ratio of $\gamma=1.5$. Furthermore, each Temba block consists of a single bidirectional Mamba layer. We follow the Mamba configurations of [50], with a state-dimension of 16. The dilation parameters are set to $\delta=\eta=b$ for Temba block $b$. For TSU, MutiTHUMOS and Charades, the input Temporal dimensions are fixed at 2500, 256 and 2000 respectively. In Appendix A, we provide the configuration details for each of the datasets. We use Adam optimizer [25] and cosine learning rate scheduler. All our models are trained on a single $24$GB RTX A5000 GPU.
Table 1: Comparison with the State-of-the-art. Bold denotes best scores, underline denotes the second best.
Method Visual #Param FLOPs TSU MTHUMOS Charades
Backbone (M) (G) (mAP) (mAP) (mAP)
Super-event [31] I3D 26 0.8 17.2 36.4 18.6
TGM [32] I3D 2 1.2 26.7 37.2 20.6
PDAN [7] I3D 6 3.2 32.7 40.2 23.7
Coarse-Fine [23] X3D 8 73.4 - - 25.1
MLAD [39] I3D 21 45.2 - 42.2 18.4
CTRN [6] I3D 11 11.3 33.5 44.0 25.3
MS-TCT [8] I3D 87 6.6 33.7 43.1 25.4
MS-Temba (ours) I3D 10 0.6 34.9 43.7 25.3
TTM [34] ViViT 89 0.8 - - 26.3
AAN [9] CLIP - - 41.3 - 32.0
MS-TCT [8] CLIP 87 6.6 40.6 41.6 31.9
MS-Temba (ours) CLIP 10 0.6 42.0 41.9 32.3
Figure 6: Illustration of the evolution of MS-Temba’s accuracy with increase in video duration. MS-Temba’s effectiveness increases considerably with increase in duration.
Figure 7: Throughput vs Video Duration. We show that as video duration increases, the throughput of MS-Temba  increases compared to the baseline.
Figure 8: Top-5 action improvements. We illustrate the top five actions where our model outperforms the baseline. These include long-duration as well as short actions.
Figure 9: Qualitative Evaluation of Action Detection Heatmap on an Example Video.
5.1 Comparison to the State-of-the-Art
In Table 1, we benchmark MS-Temba against the state-of-the-art temporal action detection methods on TSU, MultiTHUMOS and Charades. The top section of Table 1 demonstrates the performance of models using an I3D visual backbone for feature extraction, which is predominantly used in previous works [32, 7, 39, 6, 8]. In the TSU and MultiTHUMOS datasets, while other methods struggle to detect actions in the videos due to their longer duration compared to other datasets, MS-Temba achieves state-of-the-art, demonstrating its ability to learn effective long-range representations in the state space. We observe that MS-Temba achieves comparable performance with state-of-the-art [8] in Charades $88.5\%$ less number of parameters.
The bottom section of Table 1, presents the performance of Temporal models with foundation models as backbone. TTM [34] uses a ViViT backbone, pretrained on the large-scale datasets JFT [37] and Kinetics-400 [24] and finetuned on the target dataset, for feature extraction whereas the other models use the pretrained foundation model CLIP. We observe that MS-Temba  trained using CLIP features outperforms MS-Temba trained with I3D features on TSU and Charades, highlighting the fact that foundation models learn more generalizable discriminative features compared to specialized models. For a detailed discussion of the I3D vs CLIP features, see Appendix. Interestingly, we find that I3D features outperform CLIP features on the MultiTHUMOS dataset. This advantage likely stems from MultiTHUMOS’s focus on short, atomic actions, which align well with the temporal modeling capacity inherent in I3D through 3D convolutions. MS-Temba consistently outperforms the previous state-of-the-art methods across all the datasets. MS-Temba achieves a breakthrough in efficiency, utilizing $73.9\%$ less GPU memory and significantly fewer FLOPs compared to previous state-of-the-art methods. Notably, MS-Temba achieves state-of-the-art performance with 90.9% less GFLOPs when compared to MS-TCT[8].
Figures 8 and 8 present the plots of action detection accuracy and throughput versus video duration for MS-Temba and MS-TCT. These results indicate that MS-Temba, with its unique components, effectively models long temporal relationships as video duration increases. Additionally, throughput improves in the Mamba-based architecture over transformer-based architectures as video duration grows, aligning with observations of Mamba methods in other domains [50]. This efficiency makes MS-Temba well-suited for handling long sequences, such as untrimmed videos with complex temporal dependencies, which challenge transformer-based models due to their quadratic complexity.
Figure 9 provides a qualitative comparison between MS-Temba and MS-TCT, showing that MS-Temba predicts action classes more accurately, highlighting its superior performance. In Figure 8, the Top-5 improvements in action class detection over MS-TCT are presented. These improvements span both long (breakfast, use laptop) and short(stir, use oven, use cupboard) action types, as indicated by mean action duration, and occur in classes with high temporal variability, as shown by the standard deviation of action duration. Thus, MS-Temba effectively captures actions at multiple temporal scales and handles significant intra-class variation.
In Table 2, we measure MS-Temba’s ability to temporal dependencies and co-occurrence dependencies of actions on the Toyota Smarthome Untrimmed dataset, using the Action Conditioned Metrics introduced in [39]. The metrics compute Precision ($P_{AC}$), Recall ($R_{AC}$), F1-Score ($F1_{AC}$) and mAP ($mAP_{AC}$) which are conditioned on actions occurring within a temporal window $\tau$. Thus $\tau=0$ indicates co-occuring actions while $\tau>0$ suggests actions with a temporal dependency. In Table 2, we show that MS-Temba achieves competitive performance with MSTCT when $\tau=0$ and $\tau=20$, with only $\frac{1}{8}$ of the parameter size. Furthermore, it achieves state-of-the-art on $\tau=40$ which suggests that MS-Temba is effective in capturing long-range dependencies between action instances.
Table 2: Performance on Action Conditioned Metrics on TSU
Method $\tau=0$ $\tau=20$ $\tau=40$
$\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$ $\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$ $\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$
I3D [3] 12.84 3.73 5.79 13.94 24.1 6.14 9.79 21.39 25.05 6.12 9.84 20.41
I3D + MS-TCT [8] 23.26 18.15 20.4 25.74 34.19 22.67 27.26 37.12 35.49 23.86 28.53 36.35
I3D + MS-Temba (ours) 22.21 16.28 18.79 25.43 34.72 22.32 27.17 37.49 37.59 23.98 29.28 37.64
CLIP + MS-TCT [8] 26.49 21.34 23.64 31.14 40.78 30.93 35.18 43.44 42.66 32.40 36.80 43.36
CLIP + MS-Temba (ours) 25.97 22.47 24.09 29.44 42.10 30.62 35.45 43.97 43.55 32.17 37.01 44.17
5.2 Ablation Study
In this section, we ablate the different component of MS-Temba. All the ablations are conducted on the TSU and Charades datasets.
Impact of each component in MS-Temba. In this section, we aim to study the impacts of the building blocks of MS-Temba. Specifically, how the Temporal Local Module (TLM), Dilated Temporal SSM (DTS) and Temporal Mamba Fuser (TMF) are responsible for the effectiveness of MS-Temba. Table 3, demonstrates the impact of each component for action detection on Charades and TSU. We compare each model with the scenario where all the components are absent and the features extracted by the visual backbone are directly discriminated by the classification head. In Table 3, we observe that using either TLM or DTS with Temba Fuser leads to an improvement over the baseline. As mentioned in Section 4.2, TLM focuses on the local inter-token relations while DTS is responsible to capture long-range dependencies. TMF further enhances action detection accuracy by integrating temporal representations learned at different scales across blocks, as defined by the dilations in TLM and DTS. Consequently, each component of MS-Temba captures complementary information, and their combination is instrumental for effective action detection.
Table 3: Impact of Temporal Local Module (TLM), Dilated Temporal SSM (DTS) and Temporal Mamba Fuser (TMF) in MS-Temba
Temporal Local Dilated Temporal TMF TSU Charades
Module SSM (mAP) (mAP)
25.3 18.1
$\checkmark$ $\checkmark$ 37.8 31.3
$\checkmark$ $\checkmark$ 39.0 32.3
$\checkmark$ $\checkmark$ $\checkmark$ 42.0 32.3
Analysis of Temba Fuser. In Table 5, we present the strategies for fusing the outputs of each TEMBA block, to learn the multi-scale temporal representations. In the fuser, we aim to first merge the representations and then have an interaction between the merged tokens to have a unified multi-scale representation. First, following [8], we concatenate the features and have a Temporal Convolution layer for interaction among the tokens (denoted by ’Concat + Temporal Conv.’). It is observed that a linear interaction between the tokens after concatenation (denoted by ’Concat + Linear Proj.’) outperforms the former as it does not impose any specific temporal constraints leading to an effective aggregation. Furthermore, we find that aggregation through summation of the features leads to a better performance. Finally, a BiMamba interaction layer (’Sum + BiMamba Int.’) after the aggregation leads to the optimal results which demonstrates the ability of BiMamba to learn an effective state space that represents the features from each Temba block.
Table 4: Analysis of Temporal Mamba Fuser
Method TSU Ch
Concat + Temp Conv. 36.9 23.7
Concat + Lin Proj. 37.7 31.4
Sum 39.1 31.7
Sum + BiMamba Int. 42.0 32.3
Table 5: Effect of Dilation in TLM & DTS
TLM DTS TSU Ch
24.9 26.6
$\checkmark$ 41.2 30.7
$\checkmark$ 41.7 31.8
$\checkmark$ $\checkmark$ 42.0 32.3
Dilation in TLM and DTS. Here, we explore the effectiveness of dilation in the Temba block, particularly in TLM and DTS. Here, our baseline is a model which does not use dilation in either the temporal convolution in TLM or DTS. As illustrated in Table 5, we find that using dilation in either TLM or DTS results in a substantial improvement over the baseline scenario, highlighting the importance of dilation to capture multi-scale token interactions without scaling the Temporal dimension (see Appendix). Furthermore, using dilation in both TLM and DTS leads to an improvement of $21.4\%$ in Charades and $68.7\%$ increase in TSU.
5.3 Other Design Choices
In this section, we present the design choices of the Temba blocks in MS-Temba, particularly the scanning strategy of the SSM and number of Temba blocks.
Scanning Strategy. Mamba scans the input tokens sequentially. However, recent computer vision approaches [50, 26] have introduced bidirectional scanning to more effectively capture inter-token interactions in the state-space. We ablate unidirectional as well as bidirectional scanning approaches in the Mamba layers. In Table 7, we find that using a bidirectional scanning leads to optimal results owing to its ability to compute global information.
Number of Temba blocks. Table 7, we evaluate the impact of varying the number of Temba blocks in MS-Temba. Notably, the dilation rates $\delta\text{ and }\eta$ adjust in proportion to the block count, leading to larger dilation in the later blocks. It is observed that the performance peaks with three Temba blocks. Further increasing the number of blocks leads the model to learn redundant representations.
Table 6: Effect of Scanning Strategy in SSM
Scanning TSU Ch
Unidirectional 37.0 31.0
Bidirectional 42.0 32.3
Table 7: Effect of Number of Temba Blocks
#Blocks TSU Charades
1 38.6 30.8
2 39.5 31.6
3 42.0 32.3
4 25.0 31.0
(a) Comparison of Information Density of I3D and CLIP features of a sample of 15 frames averaged over 10 videos of Charades
(b) Comparison of Information Density of I3D and CLIP features of a sample of 100 frames averaged over 10 videos of TSU
Figure 10: Comparison of Information Density of I3D and CLIP features in Charades and TSU
6 Comparison of CLIP and I3D Features
The superior performance of MS-Temba in action detection, leveraging features from foundation models instead of traditional 3D ConvNets, warrants the examination of the distinctions between features extracted by a CLIP backbone and and I3D backbone — widely adopted in prior state-of-the-art approaches[8, 7, 39]. Note that CLIP is a foundational model with considerable larger number of parameters compared to I3D and it is trained on a huge dataset as outlined in [33]. I3D is pre-trained on comparatively small-scale dataset [24] and fine-tuned on the short clips of the target dataset. To facilitate this comparison, we employ information density as a metric to evaluate the representational quality of the features. In a given video $V$, the information density of a frame $v$ can be computed as:
$I(v)=\frac{1}{|V|}\sum_{v^{\prime}\in V}\textrm{sim}(v,v^{\prime})$ (13)
where sim(v,v’) denotes the similarity between features of frames $v$ and $v^{\prime}$. $|V|$ denotes the total number of frames of the video $V$. Here, we use cosine similarity to measure the similarity between frame features.
In Figure 10, we compare the information density of features extracted using a CLIP and an I3D backbone on a random sample of 15 frames for Charades (Fig. 10) and 100 frames of TSU (Fig. 10). Figure 10 shows the information density averaged over 10 videos from the respective datasets. We observe that for both Charades and TSU, owing to their superior generalization capabilities, the CLIP features possess higher average information density per frame compared to the I3D features. This enables MS-Temba to learn more effective temporal relationships from the dense CLIP features. Consequently, Mamba-inspired architectures for visual understanding have the potential to outperform transformer-based architectures when provided with features of high information density. We leave the exploration of this property of Mamba for other visual understanding tasks as future work.
7 Conclusion
In this work, we propose Multiscale Temporal Mamba (MS-Temba), leveraging the recent state-space model (SSM), Mamba, as a temporal encoder for action detection in untrimmed videos. It combines the strengths of SSMs and convolutions: convolutions model local information, while SSMs capture global dependencies. To enhance long-term relationship modeling, we introduce dilations in Mamba. This combination enables MS-Temba to learn temporal representations at multiple scales. We also utilize Mamba blocks to aggregate scale-specific information into a unified representation, which is then classified for action detection. MS-Temba achieves state-of-the-art performance on long-video datasets and matches the best-performing method on short-video datasets, using $88.5\%$ fewer parameters. Additionally, MS-Temba outperforms baseline models when leveraging visual features from foundation models. This work establishes a foundation for employing Mamba in efficient and effective long-video understanding.
Acknowledgements
We thank the members of the Genius Lab at UNC Charlotte for helpful discussions. This work is supported in part by the National Science Foundation (IIS-2245652).
Supplementary Material
8 Overview
The supplementary is categorized into the following parts:
  • Section 9: Training configurations for each Dataset
  • Section 10: Effect of Temporal Dimension Scaling
  • Section 11: Effect of the number of Mamba Layers in MS-Temba
  • Section 12: Architecture of MS-Temba
Table 8: MS-Temba Training Settings
Configuration TSU MTHUMOS Charades
Padded Temporal Length $2500$ $512$ $256$
Optimizer $AdamW$ $AdamW$ $AdamW$
Momentum $0.9$ $0.9$ $0.9$
Learning Rate (LR) $4.5e^{-4}$ $5e^{-4}$ $2.5e^{-4}$
LR Scehduler $Cosine$ $Cosine$ $Cosine$
Warmup Epochs $5$ $5$ $5$
Batch Size $1$ $4$ $2$
Total Epochs $140$ $100$ $100$
9 Training configurations for each Dataset
For all experiments, we pad the features of each input video to a fixed sequence length. Table 8 shows the configurations used for each dataset.
Table 9: Impact of scaling Temporal Dimension in MS-Temba
TMF TSU Charades
Max Pooling 30.8 31.8
Average Pooling 36.3 30.9
Strided Conv 37.8 31.2
No Scaling 42.0 32.3
Figure 11: Impact of Number of Mamba Layers in TEMBA Block
Table 10: MS-Temba architecture. The input and out feature size is following $T\times D$ format, where $T$ is the number of tokens and $D$ is the feature dimension. The hyper-parameters are defined as: $H$: linear hidden dimension, $K$: kernel size, $S$: stride, $P$: zero-padding rate, dil.: dilation rate, $D_{s}$: SSM dimension.
Block Component Learnable layer Hyper-parameters Input size Output size
Visual Backbone Projection Linear Layer H: 256 256$\times$1024 256$\times$256
Temba Block 1 TLM Temp Convolution $K$: 3, $S$: 1, $P$: 1, dil.:1 256$\times$256 256$\times$256
DTS BiMamba $D_{s}$: 256 256$\times$256 256$\times$256
Temba Block 2 TLM Temp Convolution $K$: 3, $S$: 1, $P$: 1, dil.:2 256$\times$256 256$\times$384
DTS BiMamba $D_{s}$: 384 256$\times$384 256$\times$384
Temba Block 3 TLM Temp Convolution $K$: 3, $S$: 1, $P$: 1, dil.:3 256$\times$384 256$\times$576
DTS BiMamba $D_{s}$: 576 256$\times$576 256$\times$576
Temba Fuser Projection Linear Layer H: 576 256$\times$256 256$\times$576
Linear Layer H: 576 256$\times$384 256$\times$576
Linear Layer H: 576 256$\times$576 256$\times$576
Fuser BiMamba $D_{s}$: 576 256$\times$576 256$\times$576
10 Effect of Temporal Dimension Scaling
In this section, we explore strategies for scaling the temporal dimensions in untrimmed videos to emulate the concept of hierarchical Mamba models. Similar temporal scaling technique is also employed in the state-of-the-art MS-TCT action detection model.
Table 9 illustrates the effect of scaling the temporal dimension across successive Temba blocks in MS-Temba. The temporal dimension is halved in each block, compared to its predecessor. To maintain a consistent number of tokens in the TMF, we upsample the temporal dimension to match the number of input tokens. We explore this temporal scaling through downsampling the temporal dimension using pooling methods (maxpooling, average pooling) and strided convolutions in TLM. We observe that decreasing the number of tokens through temporal downsampling leads to loss of effective information in the state space of the Temba blocks thus leading to a degradation of the model’s performance.
11 Effect of the number of BiMamba Layers in MS-Temba
Previous state-of-the-art methods, relied on stacking multiple convolutional or Transformer layers in the Temporal Encoder to expand the receptive field and capture complex patterns at higher levels of abstractions. For example, in MS-TCT [8], in each stage the Temporal Encoder consists of three Global-Local Relational blocks, i.e., three convolution and transformer layers. However, since BiMamba can capture long range dependencies effectively in its state space, in MS-Temba, it is sufficient to have a single BiMamba Layer in each Temba Block. Increasing the number of BiMamba Layers may introduce redundancies in the state-spaces of each layer, leading to a degradation of the quality of learned representations. Figure 11, shows the impact of increasing the number of BiMamba layers in Dilated Temporal SSM. For both TSU and Charades, there is a decline in the mAP of the model as the number of BiMamba layers increases.
12 Architecture of MS-Temba
Table 10 demonstrates the input and output feature size for each component of MS-Temba . First, the Visual backbone extracts features of dimension $T\times D$ (where $D=1024$ for I3D backbone and $D=768$ for CLIP backbone). These features are projected using a Linear Projection layer to $T\times 256$, which is the input to the first Temba block. Each Temba Block consists of two components: (i) Temporal Local Module (TLM) and (ii) Dilated Temporal SSM (DTS). The outputs of each Temba Block are fused in the Temba Fuser. In Table 10, we show the flow of these features in MS-Temba.

### v2

@@
 Abstract
-Action detection in real-world scenarios is particularly challenging due to densely distributed actions in hour-long untrimmed videos.
-It requires modeling both short- and long-term temporal relationships while handling significant intra-class temporal variations.
-Previous state-of-the-art (SOTA) Transformer-based architectures, though effective, are impractical for real-world deployment due to their high parameter count, GPU memory usage, and limited throughput, making them unsuitable for very long videos.
-In this work, we innovatively adapt the Mamba architecture for action detection and propose Multi-scale Temporal Mamba (MS-Temba), comprising two key components: Temporal Mamba (Temba) Blocks and the Temporal Mamba Fuser.
-Temba Blocks include the Temporal Local Module (TLM) for short-range temporal modeling and the Dilated Temporal SSM (DTS) for long-range dependencies.
-By introducing dilations, a novel concept for Mamba, TLM and DTS capture local and global features at multiple scales.
-The Temba Fuser aggregates these scale-specific features using Mamba to learn comprehensive multi-scale representations of untrimmed videos.
-MS-Temba is validated on three public datasets, outperforming SOTA methods on long videos and matching prior methods on short videos while using only one-eighth of the parameters.
+Temporal Action Detection (TAD) in untrimmed videos requires models that can efficiently (1) process long-duration videos, (2) capture temporal variations within action classes, and (3) handle dense, overlapping actions, all while remaining suitable for resource-constrained edge deployment.
+While Transformer-based methods achieve high accuracy, their quadratic complexity hinders deployment in such scenarios.
+Given the recent popularity of linear complexity Mamba-based models, leveraging them for TAD is a natural choice.
+However, naïvely adapting Mamba from language or vision tasks fails to provide an optimal solution and does not address the challenges of long, untrimmed videos.
+Therefore, we propose Multi-Scale Temporal Mamba (MS-Temba), the first Mamba-based architecture specifically designed for densely labeled TAD tasks.
+MS-Temba features Temporal Mamba Blocks (Temba Blocks), consisting of Temporal Convolutional Module (TCM) and Dilated SSM (D-SSM).
+TCM captures short-term dependencies using dilated convolutions, while D-SSM introduces a novel dilated state-space mechanism to model long-range temporal relationships effectively at each temporal scale.
+These multi-scale representations are aggregated by Scale-Aware State Fuser, which learns a unified representation for detecting densely overlapping actions.
+Experiments show that MS-Temba achieves state-of-the-art performance on long-duration videos, remains competitive on shorter segments, and reduces model complexity by 88%.
+Its efficiency and effectiveness make MS-Temba well-suited for real-world edge deployment.
 [^1] The code is available at https://github.com/thearkaprava/MS-Temba
-Figure 1: Temporal Action Detection using Mamba: On the left, a visual sample demonstrates that our model, MS-Temba, accurately localizes detected action labels in comparison to MS-TCT for temporal action detection in untrimmed videos.
-On the right, a plot of GPU memory usage versus video duration highlights the efficiency of MS-Temba in processing long untrimmed videos.
-“Temba, his arms wide.”
-Star Trek: The Next Generation episode,“Darmo”
+Figure 1: Left: Temporal Action Detection poses unique challenges, including the need for computationally efficient models capable of processing long video sequences, the presence of both short and long actions with high intra-class variance, and the complexity of densely overlapping actions.
+Right: MS-Temba, our proposed model, achieves state-of-the-art performance while being $\mathbf{9\times}$ more parameter-efficient compared to transformer-based and convolution-based approaches.
+Furthermore, its lightweight architecture enables seamless deployment on edge devices, underscoring its efficiency and practicality for real-world applications.
 1 Introduction
-Learning representations in long, untrimmed videos is challenging due to complex, long-term dependencies that advanced neural networks struggle to optimize end-to-end.
-One such task, action detection, involves precisely localizing temporal boundaries of actions in untrimmed videos, with applications in areas like patient monitoring and video surveillance.
-Real-world deployment of temporal action detection models requires efficiency in capturing actions and their boundaries, yet real-world videos differ significantly from sports datasets [22, 2], where a few, non-overlapping actions with clear motion patterns are present.
-Real-world videos often contain densely distributed actions over time with subtle motion patterns, further complicated by large intra-class variance (e.g., the same action occurring with varying durations), the need for long-term relation modeling, and simultaneous actions.
-This paper focuses on developing an efficient model for action detection in hour-long videos, addressing these real-world challenges.
-To learn temporal representations in untrimmed videos, early methods employed proposal-based approaches [43], which, while effective for videos with sparse action distributions, are resource-intensive due to the need to generate numerous proposals.
-Consequently, temporal convolutional methods [14, 31, 32] became popular for action detection, although their local receptive fields limit the capture of long-term dependencies.
-With the success of self-attention mechanisms, transformer-based methods [39, 8] were introduced to learn global representations across time frames.
-Despite their effectiveness, these transformer models impose high computational costs in terms of model parameters, memory usage, and throughput, making them challenging to deploy on edge devices for real-world applications.
-In recent research, Mamba [17, 26, 50] is gaining attention for its efficiency in handling long sequences with linear complexity, making it well-suited for action detection in untrimmed videos.
-While Mamba has seen exploration in the vision domain, its application to action detection in untrimmed videos remains underexplored, motivating us to develop a Mamba-based architecture tailored for this purpose.
-To address this, we propose Multi-scale Temporal Mamba (MS-Temba), an architecture designed for efficient action detection (see Figure 1).
-MS-Temba incorporates two key components: 1) Temporal Mamba Blocks (Temba blocks) and 2) the Temporal Mamba Fuser (Temba fuser).
-The Temba Blocks are designed to learn temporal representations at multiple scales, achieved through two modules: the Temporal Local Module (TLM) and the Dilated Temporal State Space Module (DTS).
-TLM captures local temporal representations, while DTS learns long-range relationships using state space models enhanced with dilations.
-Although dilated convolutions are widely used, this work is the first to incorporate dilations into SSMs, allowing DTS to skip immediate neighbors and capture relationships across different temporal scales.
-This approach enhances Temba Blocks’ ability to model long-term temporal relations effectively.
-Additionally, the Temba Fuser aggregates scale-specific temporal information from each Temba Block, learning a multi-scale representation.
-This fusion, performed via Mamba, is computationally lighter than dense layers.
-We assess MS-Temba on three publicly available datasets with densely labeled actions, finding that it performs on-par with previous state-of-the-art method on short-duration videos and surpasses it on long-duration videos, using only one-eighth the parameters.
-Our study reveals that MS-Temba outperforms transformer-based models when both use a foundational visual backbone, highlighting Mamba’s potential not only for long sequences but also for feature inputs from foundation models [33].
+[…]
+To learn temporal representations in untrimmed videos, early methods relied on proposal-based approaches [53], which, while effective for sparsely distributed actions, are computationally expensive due to dense proposal generation.
+To improve efficiency, Temporal Convolutional Networks (TCNs) [14, 39] have been explored, but their local receptive fields limit their ability to capture long-range dependencies.
+Recently, Transformer-based methods [49, 8] have achieved strong performance in TAD by leveraging global self-attention.
+However, their quadratic complexity leads to excessive memory consumption, making them impractical for long video sequences and edge deployment.
+Given the growing need for efficient long-sequence modeling, state-space models (SSMs) such as Mamba [18] have emerged, offering linear complexity while effectively capturing long-term dependencies.
+Although Mamba has demonstrated promising direction in vision tasks [18, 60, 27, 4], its application to TAD in untrimmed videos remains underexplored, motivating us to develop a Mamba-based architecture tailored for this purpose.
+A naïve approach to adapting Mamba for TAD is to directly apply 1D Mamba to temporal segment representations, which are pooled frame-level features from a visual backbone.
+However, this approach sacrifices local information by compressing all context into a single state [18].
+Additionally, Mamba sequentially processes tokens and relies on memorization, limiting its ability to model long-range temporal dependencies compared to Transformer counterparts.
+These limitations hinder its effectiveness in detecting actions in untrimmed videos.
+Therefore, we propose Multi-Scale Temporal Mamba (MS-Temba), a novel Mamba-based architecture designed for capturing temporal representations in videos.
+MS-Temba comprises two key components: (1) Temporal Mamba Blocks (Temba Blocks) and (2) Scale-Aware State Fuser (SSF), each tailored to address specific challenges in TAD.
+Temba Blocks capture both fine-grained and long-term dependencies by learning representations at multiple temporal scales.
+Each Temba Block consists of two modules: Temporal Convolutional Module (TCM) and Dilated SSM (D-SSM).
+TCM captures short-term dependencies using dilated convolutions, while D-SSM extends Mamba’s capability by incorporating dilations in the state-space update mechanism.
+The combination of TCM and D-SSM enables Mamba to learn discriminative temporal representations, with dilation parameters facilitating multi-scale temporal learning.
+To handle dense, overlapping actions spanning both short and long durations, effective aggregation of scale-aware temporal representations is essential.
+This is achieved through our proposed Scale-Aware State Fuser, which learns a unified representation across multiple temporal scales using SSMs.
+Our experimental analysis shows that MS-Temba outperforms state-of-the-art performance on long-duration videos while maintaining comparable accuracy to Transformer-based models on short-duration videos, using only one-eighth of the parameters.
+Additionally, MS-Temba outperforms Transformer-based models when both utilize a foundational visual backbone like CLIP [40], demonstrating Mamba’s superior effectiveness in high-information-density settings.
+While “local-global” representation learning and “multi-scale temporal learning” have been studied in prior work [14, 8, 35, 13, 28], MS-Temba is the first attempt to learn such representations leveraging SSMs for TAD.
+This paper presents a systematic approach to adapting the Mamba architecture for TAD, yielding key findings.
+Our carefully designed strategy enables MS-Temba to be trained directly on edge devices, providing advantages in privacy, personalization, efficiency, and real-time adaptability, making it well-suited for wearables and autonomous systems (see Figure 1).
 To summarize the main contributions:
   1. 1.
-    We introduce a novel Mamba-based architecture Multi-scale Temporal Mamba (MS-Temba), tailored for densely labeled Temporal Action Detection
+    We introduce Multi-Scale Temporal Mamba (MS-Temba), the first Mamba-based architecture tailored for densely labeled Temporal Action Detection.
   2. 2.
-    MS-Temba explores different temporal scales through our novel components in Temba block, especially dilated Temporal SSMs, and aggregates them to learn multiscale representation through the Temporal Mamba Fuser
+    We propose Temba Blocks, consisting of TCM and D-SSM.
+TCM captures short-term action dependencies using convolutions, while our novel D-SSM module introduces dilations in SSMs to model long-range temporal contexts.
+Additionally, we introduce Scale-Aware State Fuser, which efficiently aggregates multi-scale temporal representations, enabling dense action localization.
   3. 3.
-    This carefully crafted design of MS-Temba  achieves performance on par with Transformer-based architectures on short videos and outperforms the previous state-of-the-art on long-duration videos with $88\%$ less parameters.
-Figure 1 illustrates a visual sample of action detection using MS-Temba compared to previous state-of-the-art methods.
-MS-Temba efficiently processes hour-long videos with low GPU memory usage and fewer parameters, maintaining high accuracy.
-This motivates the suitability of the Mamba architecture for temporal action detection in untrimmed videos.
+    MS-Temba surpasses prior state-of-the-art methods on long-duration videos while maintaining competitive performance on shorter videos, with 88% fewer parameters.
 2 Related Work
+Figure 2: Overall Architecture of Multiscale Temporal Mamba (MS-Temba) for action detection.
+MS-Temba is composed of a frozen pretrained Visual Backbone, Temporal Mamba (Temba) Blocks for learning local-global representations at multiplle scale spaces.
+The Scale-Aware State Fuser employs an SSM for effectively fusing the multi-scale features which is projected by the Classification Head for dense action detection.
+Temporal Action Detection.
 Understanding actions in long temporal sequences has become a key task in computer vision [17, 30, 54, 11, 10, 58].
@@
 However, temporal action detection with densely labeled action distribution offers greater challenges due to concurrent actions, where proposal-based methods struggle with the combinatorial explosion in proposal generation.
-This led to the adoption of temporal convolution-based models [31, 14] for processing long videos.
+This led to the adoption of temporal convolution-based models [38, 14, 31, 32] for processing long videos.
 Convolutional methods like PDAN [7] and TGM [39] use specialized kernels designed to model long-term dependencies and capture composite actions.
@@
 This limitation is significant for densely labeled Temporal Action Detection, where long-term actions are common.
-In response, transformer-based architectures [41, 12, 40, 29, 13], which have proven effective in computer vision, have been adapted for action detection [39, 8, 34].
+In response, transformer-based architectures [51, 12, 50, 35, 13], which have proven effective in computer vision, have been adapted for action detection [49, 8, 43, 47, 16, 42].
 These architectures can model both short- and long-term dependencies in videos.
@@
 These transformer based approaches enhances multi-scale representation learning, but the effectiveness comes at a high computational cost.
-The linear complexity of state space models (Mamba [17]) makes it an ideal test bed for processing long untrimmed videos.
+The linear complexity of state space models [18] makes it an ideal test bed for processing long untrimmed videos.
+State Space Models (SSMs).
 In recent times, Structured State Space Model (S4) [19] has gained popularity in effectively processing long sequence lengths in linear complexity.
@@
 To the best of our knowledge, MS-Temba is the first to leverage Mamba for the task of temporal action detection in densely labeled untrimmed videos.
-3 Preliminaries
-State space models map continuous input sequences $x(t)\in\mathbb{R}$ to continuous outputs $y(t)\in\mathbb{R}$ using a hidden state $h(t)\in\mathbb{R}^{n}$.
-For the above mapping and updating of the hidden state, SSMs use a transition state matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ projection matrices $\mathbf{B}\in\mathbb{R}^{n\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times n}$.
-The update equations are given by
-$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{A}x(t)$ (1)
-$y(t)$ $=\mathbf{C}h(t)$
-Previous works  [18, 17, 50], use discrete versions of the above system using a discretization parameter $\mathbf{\Delta}$ known as the step size.
-Following is an example of a discretization transformation using $\mathbf{\Delta}$, known as zero-order hold (ZOH):
-$\overline{\mathbf{A}}$ $=\exp(\mathbf{\Delta}\mathbf{A})$ (2)
-$\overline{\mathbf{B}}$ $=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp(\mathbf{\Delta}\mathbf{A})-\mathbf{I})\mathbf{\Delta}\mathbf{B}$
-This discretization of $\mathbf{A}$, $\mathbf{B}$ to $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ leads to the following updated state-space equations:
-$h_{t}$ $=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t}$ (3)
-$y_{t}$ $=\mathbf{C}h_{t}$
-Standard SSMs are typically designed to process 1D inputs in a linear sequence.
-However, [50, 26] introduce bidirectional scanning of tokens (illustrated in Figure 3), which are more effective for vision tasks.
-In this approach, the SSM operates on the input sequence $v$ and it’s reversed counterpart $v^{\prime}$, fusing the projected outputs for robust bidirectional sequence modeling.
-For brevity, we call it BiMamba for the remainder of the paper.
-This enhancement enables more effective representation learning, especially in domains requiring complex temporal dependencies.
-In this work, we extend the Mamba based architectures for understanding long untrimmed videos.
-4 Multi-scale Temporal Mamba
-Figure 2: Overall Architecture of Multiscale Temporal Mamba (MS-Temba) for action detection.
-MS-Temba is composed of a Visual Backbone, a Temporal Encoder consisting of Temporal Mamba (Temba) Blocks, a Temporal Mamba Fuser and a Classification Head.
+3 Multi-scale Temporal Mamba (MS-Temba)
 In Temporal Action Detection, we aim to classify multiple activities at each time-step of an untrimmed video.
 Formally, given a video of $T$ temporal segments, we associate an action label $y^{c}_{t}$ for each time-step $t$, where $c$ denotes the action classes, with $c\in\{1,...,C\}$.
-Here, we introduce Multi-Scale Temporal Mamba (MS-Temba), which is capable of detecting fine-grained as well as composite actions efficiently in long untrimmed videos.
-As illustrated in Figure 2, MS-Temba  consists of: i) a Visual Backbone to extract discriminative features from untrimmed videos, ii) Temba blocks which captures temporal information at different scales, iii) Temba Fuser to combine the learned representations across different scales, and iv) a Classification Head to predict multiple actions within each temporal segment of the input video.
-In the following sections, we detail each of these components of MS-Temba.
+Here, we introduce Multi-Scale Temporal Mamba (MS-Temba), which is capable of detecting long as well as short actions efficiently in long untrimmed videos.
+As illustrated in Figure 2, MS-Temba  consists of: i) a Visual Backbone to extract discriminative features from untrimmed videos, ii) Temba blocks which capture temporal information at different scales, iii) Scale-Aware State Fuser to combine the learned temporal representations across different scales, and iv) a Classification Head to predict multiple actions within each temporal segment of the input video.
+In the following sections, we detail each of these components.
 3.1 Visual Backbone
@@
 The segment-wise representations are then stacked to form an input feature vector $v\in\mathbb{R}^{T\times D}$, where D is the feature dimension and T is the total temporal sequence length.
-Finally, $v$ is projected to $z_{0}\in\mathbb{R}^{T\times D_{0}}$ to obtain a latent representation of the input, which serves as input tokens to our proposed Mamba-based temporal encoder.
+Finally, $v$ is projected to $z_{0}\in\mathbb{R}^{T\times D_{0}}$ to obtain a latent representation of the input, which serves as input tokens to our proposed Temba Blocks.
 Figure 3: Temba Block.
-A single Temba block consists of a Temporal Local Module (TLM) and a Dilated Temporal SSM.
-The TLM is responsible for learning local feature while the DTS learns long-range temporal dependencies.
+A single Temba block consists of a Temporal Convolutional Module (TCM) and Dilated SSM (D-SSM).
+The TCM is responsible for capturing local information while the D-SSM learns long-range temporal dependencies.
 $\Phi$ denotes a non-parameterized operation to facilitate dilation in the SSM.
@@
 To this end, we introduce Temporal Mamba (Temba) Blocks which consists of two primary modules.
-i) Temporal Local Module to capture local temporal relations between tokens and ii) Dilated Temporal SSM to capture long-range temporal dependencies within videos.
-Temporal Local Module (TLM) forms an essential component of Temba blocks, specifically designed to introduce local interactions between tokens in neighboring temporal region.
+i) Temporal Convolutional Module (TCM) to capture local relations between tokens and ii) Dilated SSM (D-SSM) to capture long-range temporal dependencies in videos.
+Temporal Convolutional Module (TCM) forms an essential component of Temba blocks, specifically designed to introduce local interactions between tokens in neighboring temporal region.
 In order to capture multiscale temporal representations it is important to process the tokens locally at different neighborhoods.
-Therefore, TLM uses a dilated temporal convolution layer with a kernel size of $k=3$, stride of $1$ and dilation of $\delta$.
+Therefore, TCM uses a dilated temporal convolution layer with a kernel size of $k=3$, stride of $1$ and dilation of $\delta$.
 We gradually increase $\delta$ in subsequent blocks.
@@
 $z_{b}=W_{b}z_{b-1}+\beta_{b},\quad z_{b}\in\mathbb{R}^{B\times T\times D_{b}}$ (1)
-where $W_{b}$ is the 1D convolutional kernel of the $b^{th}$ Temba Block and $\beta_{b}$ is the bias, $D_{b}$ is the feature dimension computed as $D_{b}=\gamma D_{b-1}$.
-Thus, the TLM progressively increases the feature dimension by a factor of $\gamma$ in each successive block, enabling the model to capture complex patterns effectively as its receptive field increases.
+where $B$ is the batch size, $W_{b}$ is the 1D convolutional kernel of the $b^{th}$ Temba Block and $\beta_{b}$ is the bias, $D_{b}$ is the feature dimension computed as $D_{b}=\gamma D_{b-1}$.
+Thus, the TCM progressively increases the feature dimension by a factor of $\gamma$ in each successive block, enabling the model to capture complex patterns effectively as its receptive field increases.
 The receptive field for a kernel of size 3 and dilation $\delta$ is given by,
 $ReceptiveField(\delta)=2^{\delta}+1$ (2)
-Functionalities of the TLM can also be achieved using MLPs or more complex networks for local interaction.
-Consequently, the TLM introduces a strong inductive bias by mapping token representations across multiple temporal scales, subsequently enabling the temporal SSM to capture global token interactions effectively.
-Figure 4: Scanning Strategy.
-An illustration of scanning strategy in Dilated Temporal SSM (DTS).
-Note this diagram shows only the forward scanning in the second Temba Block.
-MS-Temba uses bidirectional scanning.
-Dilated Temporal SSM.
-(DTS) While the TLM captures interactions in a local neighborhood, it is not sufficient for learning long-range temporal relationships among actions which is crucial for Temporal Action Detection, as discussed in Section 1.
+Functionalities of the TCM can also be achieved using MLPs or more complex networks for local interaction.
+Consequently, the TCM introduces a strong inductive bias by mapping token representations across multiple temporal scales, subsequently enabling the temporal SSM to capture global token interactions effectively.
+Figure 4: Scanning Strategy. An illustration of scanning strategy in Dilated SSM (D-SSM). Note this diagram shows only the forward scanning in the Temba Blocks. MS-Temba uses bidirectional scanning.
+Dilated SSM (D-SSM).
+While the TCM captures interactions in a local neighborhood, it is not sufficient for learning long-range temporal relationships among actions which is crucial for Temporal Action Detection, as discussed in Section 1.
 These global interactions are effective in Transformer-based architectures, but they come at a high computational cost and are infeasible for hour-long videos.
@@
 These are scanned separately by a bidirectional SSM (BiMamba), illustrated in Figure 4, and the learned representations are merged together in the respective initial positions (as in $T$) to achieve a global interaction among the tokens in the SSM.
-This approach, dubbed Dilated Temporal SSM, is formally described as follows:
+This approach, dubbed Dilated SSM, is formally described as follows:
 $z_{b}$ $=\mathit{\Phi}_{\eta}(z_{b}),\quad z_{b}\in\mathbb{R}^{\eta B\times T/\eta\times D}$ (3)
-$z_{b}$ $=BiMamba(z_{b}),\quad z_{b}\in\mathbb{R}^{\eta B\times T/\eta\times D}$ (7)
+$z_{b}$ $=\text{BiMamba}(z_{b}),\quad z_{b}\in\mathbb{R}^{\eta B\times T/\eta\times D}$ (4)
 $z_{b}$ $=\mathit{\Phi}^{-1}_{\eta}(z_{b}),\quad z_{b}\in\mathbb{R}^{B\times T\times D}$ (5)
-The combination of TLM with dilated token scanning enables the Temba blocks to refine state representations across layers at various temporal scales, defined by the two dilation parameters, $\delta$ and $\eta$.
-Figure 5: Temporal Mamba Fuser (TMF) is composed of projection layers to aggregate features learnt at different Temba Blocks.
-After these features are fused, BiMamba is responsible for interaction between the fused features.
-4.3 Temporal Mamba Fuser
-The Temoral Mamba Fuser (Temba Fuser), illustrated in Figure 5, plays a crucial role in aggregating temporal features learned across different scales through Temba blocks, enabling the learning of multi-scale temporal representations.
+The combination of TCM with dilated token scanning enables the Temba blocks to refine state representations across layers at various temporal scales, defined by the two dilation parameters, $\delta$ and $\eta$.
+Figure 5: Scale-Aware State Fuser (SSF) is composed of projection layers to aggregate features learnt at different Temba Blocks.
+After these features are fused, BiMamba is responsible for learning a discriminative scale-aware state space.
+3.3 Scale-Aware State Fuser (SSF)
+The Scale-Aware State Fuser (SSF), illustrated in Figure 5, plays a crucial role in aggregating temporal features learned across different scales through Temba blocks, enabling the learning of multi-scale temporal representations.
 It first applies a linear projection to the output of each block standardizing the feature dimensionality to match that of the final block ($E$).
 The resulting features are fused to form a unified temporal representation.
-Finally, we leverage a BiMamba to facilitate interactions across the fused representation from various blocks, each capturing distinct temporal patterns, as follows:
+Finally, we leverage a BiMamba to learn state-space representation from various Temba blocks, each capturing distinct scale-aware temporal patterns, as follows:
 $z_{f}$ $=\sum_{b}W_{f}z_{b}+\beta_{f},\quad z_{f}\in\mathbb{R}^{B\times T\times E}$ (6)
-$y^{\prime}$ $=BiMamba(z_{f}),\quad y^{\prime}\in\mathbb{R}^{B\times T\times E}$ (10)
+$y^{\prime}$ $=\text{BiMamba}(z_{f}),\quad y^{\prime}\in\mathbb{R}^{B\times T\times E}$ (7)
 Here, $W_{f}$ and $\beta_{f}$ denote the weight matrix and bias of the linear projector.
 3.4 Classification Head
-The Classification Head receives the fused features from the Temba Fuser post interaction in the BiMamba.
-This component is designed as a linear layer that projects the high-dimensional temporal features into the space of multi-label action classes.
-$\tilde{y}=W_{h}y^{\prime}+B_{h},\quad\tilde{y}s\in\mathbb{R}^{B\times T\times C}$ (11)
+The Classification Head receives the scale-aware fused features from the SSF.
+This component is a linear layer that projects the high-dimensional temporal features into the space of multi-label action classes.
+$\tilde{y}=W_{h}y^{\prime}+B_{h},\quad\tilde{y}\in\mathbb{R}^{B\times T\times C}$ (8)
 Here, $y^{\prime}\in\mathbb{R}^{B\times T\times E}$ is the representation obtained from Temba Fuser, $W_{h}$ and $B_{h}$ are the parameters of the classification head projector and C is the number of action classes.
@@
 It contains 66500 annotations of daily activities with an average of 6.8 action instances per video.
-We evaluate our methods on these datasets using the ”action localization” protocol described in [5][36], and [46], respectively, to detect frame-level actions and compare them with the temporal annotations.
+We evaluate our methods on these datasets using the ‘action localization’ protocol described in [5][45], and [56], respectively, to detect frame-level actions and compare them with the temporal annotations.
+Figure 6: Model performance vs Video Duration. MS-Temba’s effectiveness increases considerably with increase in duration.
+Figure 7: Throughput vs Video Duration. We show that as video duration increases, the throughput of MS-Temba  increases compared to the baseline.
+Figure 8: Top-5 improving actions. Top five actions where our model outperforms the baseline. These include long & short actions with intra-class variance.
 Implementation Details.
@@
 For TSU, MutiTHUMOS and Charades, the input Temporal dimensions are fixed at 2500, 256 and 2000 respectively.
-In Appendix A, we provide the configuration details for each of the datasets.
+In the Supplementary material, we provide the configuration details for each of the datasets.
 We use Adam optimizer [26] and cosine learning rate scheduler.
 All our models are trained on a single $24$GB RTX A5000 GPU.
+Additionally, for experiments on edge-device, we deploy our model on an NVIDIA Jetson Nano (8GB) with a 15W power configuration.
 Table 1: Comparison with the State-of-the-art. Bold denotes best scores, underline denotes the second best.
@@
 MLAD [49] I3D 21 45.2 - 42.2 18.4
+PointTAD [47] I3D - - - 41.2 22.1
+HAAN [16] I3D - - - 41.7 25.2
 CTRN [6] I3D 11 11.3 33.5 44.0 25.3
 MS-TCT [8] I3D 87 6.6 33.7 43.1 25.4
+DualDETR [61] I3D 21 5.5 34.8 45.5 23.2
 MS-Temba (ours) I3D 10 0.6 34.9 43.7 25.3
@@
 MS-Temba (ours) CLIP 10 0.6 42.0 41.9 32.3
-Figure 6: Illustration of the evolution of MS-Temba’s accuracy with increase in video duration. MS-Temba’s effectiveness increases considerably with increase in duration.
-Figure 7: Throughput vs Video Duration. We show that as video duration increases, the throughput of MS-Temba  increases compared to the baseline.
-Figure 8: Top-5 action improvements. We illustrate the top five actions where our model outperforms the baseline. These include long-duration as well as short actions.
-Figure 9: Qualitative Evaluation of Action Detection Heatmap on an Example Video.
 4.1 Comparison to the State-of-the-Art
@@
 The top section of Table 1 demonstrates the performance of models using an I3D visual backbone for feature extraction, which is predominantly used in previous works [39, 7, 49, 6, 8].
-In the TSU and MultiTHUMOS datasets, while other methods struggle to detect actions in the videos due to their longer duration compared to other datasets, MS-Temba achieves state-of-the-art, demonstrating its ability to learn effective long-range representations in the state space.
-We observe that MS-Temba achieves comparable performance with state-of-the-art [8] in Charades $88.5\%$ less number of parameters.
-The bottom section of Table 1, presents the performance of Temporal models with foundation models as backbone.
-TTM [34] uses a ViViT backbone, pretrained on the large-scale datasets JFT [37] and Kinetics-400 [24] and finetuned on the target dataset, for feature extraction whereas the other models use the pretrained foundation model CLIP.
-We observe that MS-Temba  trained using CLIP features outperforms MS-Temba trained with I3D features on TSU and Charades, highlighting the fact that foundation models learn more generalizable discriminative features compared to specialized models.
-For a detailed discussion of the I3D vs CLIP features, see Appendix.
-Interestingly, we find that I3D features outperform CLIP features on the MultiTHUMOS dataset.
-This advantage likely stems from MultiTHUMOS’s focus on short, atomic actions, which align well with the temporal modeling capacity inherent in I3D through 3D convolutions.
-MS-Temba consistently outperforms the previous state-of-the-art methods across all the datasets.
-MS-Temba achieves a breakthrough in efficiency, utilizing $73.9\%$ less GPU memory and significantly fewer FLOPs compared to previous state-of-the-art methods.
-Notably, MS-Temba achieves state-of-the-art performance with 90.9% less GFLOPs when compared to MS-TCT[8].
+In the TSU dataset, while other methods struggle to detect actions in the videos due to their longer duration compared to other datasets, MS-Temba outperforms the state-of-the-art methods, demonstrating its ability to learn effective long-range representations in the state space.
+We observe that MS-Temba achieves comparable performance with state-of-the-art [8] in Charades and MultiTHUMOS datasets $88.5\%$ less number of parameters.
+[…]
+Finding 1: Mamba has superior effectiveness in high-information-density settings typical in features extracted from foundation models.
+Table 2: Performance on Action Conditioned Metrics on TSU.
+$P_{AC}$ Action Conditioned Precision; $R_{AC}$ Action Conditioned Recall; $F1_{AC}$ Action Conditioned F1-score; $mAP_{AC}$ Action Conditioned Mean Average Precision.
+$\tau$ indicates the temporal window size.
+Method $\tau=0$ $\tau=20$ $\tau=40$
+$\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$ $\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$ $\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$
+I3D [3] 12.84 3.73 5.79 13.94 24.10 6.14 9.79 21.39 25.05 6.12 9.84 20.41
+I3D + MS-TCT [8] 23.26 18.15 20.40 25.74 34.19 22.67 27.26 37.12 35.49 23.86 28.53 36.35
+I3D + MS-Temba (ours) 22.21 16.28 18.79 25.43 34.72 22.32 27.17 37.49 37.59 23.98 29.28 37.64
+CLIP + MS-TCT [8] 26.49 21.34 23.64 31.14 40.78 30.93 35.18 43.44 42.66 32.40 36.80 43.36
+CLIP + MS-Temba (ours) 25.97 22.47 24.09 29.44 42.10 30.62 35.45 43.97 43.55 32.17 37.01 44.17
 Figures 8 and 8 present the plots of action detection accuracy and throughput versus video duration for MS-Temba and MS-TCT.
-These results indicate that MS-Temba, with its unique components, effectively models long temporal relationships as video duration increases.
-Additionally, throughput improves in the Mamba-based architecture over transformer-based architectures as video duration grows, aligning with observations of Mamba methods in other domains [50].
-This efficiency makes MS-Temba well-suited for handling long sequences, such as untrimmed videos with complex temporal dependencies, which challenge transformer-based models due to their quadratic complexity.
-Figure 9 provides a qualitative comparison between MS-Temba and MS-TCT, showing that MS-Temba predicts action classes more accurately, highlighting its superior performance.
-In Figure 8, the Top-5 improvements in action class detection over MS-TCT are presented.
-These improvements span both long (breakfast, use laptop) and short(stir, use oven, use cupboard) action types, as indicated by mean action duration, and occur in classes with high temporal variability, as shown by the standard deviation of action duration.
-Thus, MS-Temba effectively captures actions at multiple temporal scales and handles significant intra-class variation.
-In Table 2, we measure MS-Temba’s ability to temporal dependencies and co-occurrence dependencies of actions on the Toyota Smarthome Untrimmed dataset, using the Action Conditioned Metrics introduced in [39].
+These results indicate that MS-Temba, with its unique components, effectively models long temporal relationships as the duration of the video increases.
+Furthermore, throughput improves in the Mamba-based architecture over transformer-based architectures as the video duration increases, aligning with observations of the Mamba methods in other domains [60].
+This efficiency makes MS-Temba well-suited for handling long sequences, such as untrimmed videos, spanning $>$20 mins, with complex temporal dependencies, which challenge transformer-based models due to their quadratic complexity.
+[…]
+In Table 2, we measure MS-Temba’s ability to temporal dependencies and co-occurrence dependencies of actions on the TSU dataset, using the Action Conditioned Metrics introduced in [49].
 The metrics compute Precision ($P_{AC}$), Recall ($R_{AC}$), F1-Score ($F1_{AC}$) and mAP ($mAP_{AC}$) which are conditioned on actions occurring within a temporal window $\tau$.
@@
 In Table 2, we show that MS-Temba achieves competitive performance with MSTCT when $\tau=0$ and $\tau=20$, with only $\frac{1}{8}$ of the parameter size.
-Furthermore, it achieves state-of-the-art on $\tau=40$ which suggests that MS-Temba is effective in capturing long-range dependencies between action instances.
-Table 2: Performance on Action Conditioned Metrics on TSU
-Method $\tau=0$ $\tau=20$ $\tau=40$
-$\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$ $\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$ $\mathbf{P}_{AC}$ $\mathbf{R}_{AC}$ $\mathbf{F1}_{AC}$ $\mathbf{mAP}_{AC}$
-I3D [3] 12.84 3.73 5.79 13.94 24.1 6.14 9.79 21.39 25.05 6.12 9.84 20.41
-I3D + MS-TCT [8] 23.26 18.15 20.4 25.74 34.19 22.67 27.26 37.12 35.49 23.86 28.53 36.35
-I3D + MS-Temba (ours) 22.21 16.28 18.79 25.43 34.72 22.32 27.17 37.49 37.59 23.98 29.28 37.64
-CLIP + MS-TCT [8] 26.49 21.34 23.64 31.14 40.78 30.93 35.18 43.44 42.66 32.40 36.80 43.36
-CLIP + MS-Temba (ours) 25.97 22.47 24.09 29.44 42.10 30.62 35.45 43.97 43.55 32.17 37.01 44.17
+Furthermore, it achieves state-of-the-art on $\tau=40$ which again confirms that MS-Temba is effective in capturing long-range dependencies between action instances.
+Finding 2: In TAD, Mamba is more effective than Transformers in long duration videos compared to shorter clips.
+Table 3: Performance on Edge Device. Comparison of latency, memory and power consumption when MS-Temba is deployed on an edge device.
+Method Training Memory Power TSU MTHUMOS Average
+Latency (s) (MB) (Watt) (mAP) (mAP) (mAP)
+Linear Model 0.4 0.05 5.6 23.6 36.3 30.0
+MS-TCT [8] 10.6 4 11.2 34.7 11.2 23.0
+MS-Temba 4.2 1 8 39.7 38.9 39.3
+[…]
+Table 4: Impact of TCM, D-SSM  and SSF in MS-Temba TCM D-SSM SSF TSU Ch 25.3 18.1 $\checkmark$ $\checkmark$ 37.8 31.3 $\checkmark$ $\checkmark$ 39.0 32.3 $\checkmark$ $\checkmark$ $\checkmark$ 42.0 32.3 Table 5: Analysis of Scale-Aware State Fuser Method TSU Ch Concat + Temp Conv.
+36.9 23.7 Concat + Lin Proj.
+37.7 31.4 Sum 39.1 31.7 Sum + BiMamba Int.
+42.0 32.3
 4.2 Ablation Study
@@
 In this section, we aim to study the impacts of the building blocks of MS-Temba.
-Specifically, how the Temporal Local Module (TLM), Dilated Temporal SSM (DTS) and Temporal Mamba Fuser (TMF) are responsible for the effectiveness of MS-Temba.
+Specifically, how the Temporal Convolutional Module (TCM), Dilated SSM (D-SSM) and Scale-Aware State Fuser (SSF) are responsible for the effectiveness of MS-Temba.
 Table 5, demonstrates the impact of each component for action detection on Charades and TSU.
 We compare each model with the scenario where all the components are absent and the features extracted by the visual backbone are directly discriminated by the classification head.
-In Table 3, we observe that using either TLM or DTS with Temba Fuser leads to an improvement over the baseline.
-As mentioned in Section 4.2, TLM focuses on the local inter-token relations while DTS is responsible to capture long-range dependencies.
-TMF further enhances action detection accuracy by integrating temporal representations learned at different scales across blocks, as defined by the dilations in TLM and DTS.
+In Table 5, we observe that using either TCM or D-SSM with SSF leads to an improvement over the baseline.
+As mentioned in Section 3.2, TCM focuses on the local inter-token relations while D-SSM is responsible to capture long-range dependencies.
+SSF further enhances action detection accuracy by integrating temporal representations learned at different scales across blocks, as defined by the dilations in TCM and D-SSM.
 Consequently, each component of MS-Temba captures complementary information, and their combination is instrumental for effective action detection.
-Table 3: Impact of Temporal Local Module (TLM), Dilated Temporal SSM (DTS) and Temporal Mamba Fuser (TMF) in MS-Temba
-Temporal Local Dilated Temporal TMF TSU Charades
-Module SSM (mAP) (mAP)
-25.3 18.1
-$\checkmark$ $\checkmark$ 37.8 31.3
-$\checkmark$ $\checkmark$ 39.0 32.3
-$\checkmark$ $\checkmark$ $\checkmark$ 42.0 32.3
-Analysis of Temba Fuser.
-In Table 5, we present the strategies for fusing the outputs of each TEMBA block, to learn the multi-scale temporal representations.
-In the fuser, we aim to first merge the representations and then have an interaction between the merged tokens to have a unified multi-scale representation.
-First, following [8], we concatenate the features and have a Temporal Convolution layer for interaction among the tokens (denoted by ’Concat + Temporal Conv.’).
-It is observed that a linear interaction between the tokens after concatenation (denoted by ’Concat + Linear Proj.’) outperforms the former as it does not impose any specific temporal constraints leading to an effective aggregation.
-Furthermore, we find that aggregation through summation of the features leads to a better performance.
-Finally, a BiMamba interaction layer (’Sum + BiMamba Int.’) after the aggregation leads to the optimal results which demonstrates the ability of BiMamba to learn an effective state space that represents the features from each Temba block.
-Table 4: Analysis of Temporal Mamba Fuser
-Method TSU Ch
-Concat + Temp Conv. 36.9 23.7
-Concat + Lin Proj. 37.7 31.4
-Sum 39.1 31.7
-Sum + BiMamba Int. 42.0 32.3
-Table 5: Effect of Dilation in TLM & DTS
-TLM DTS TSU Ch
+Table 6: Effect of Dilation in TCM & D-SSM
+TCM D-SSM TSU Ch
 24.9 26.6
@@
 $\checkmark$ $\checkmark$ 42.0 32.3
-Dilation in TLM and DTS.
-Here, we explore the effectiveness of dilation in the Temba block, particularly in TLM and DTS.
-Here, our baseline is a model which does not use dilation in either the temporal convolution in TLM or DTS.
-As illustrated in Table 5, we find that using dilation in either TLM or DTS results in a substantial improvement over the baseline scenario, highlighting the importance of dilation to capture multi-scale token interactions without scaling the Temporal dimension (see Appendix).
-Furthermore, using dilation in both TLM and DTS leads to an improvement of $21.4\%$ in Charades and $68.7\%$ increase in TSU.
+Table 7: Effect of Scanning Strategy in SSM
+Scanning TSU Ch
+Unidirectional 37.0 31.0
+Bidirectional 42.0 32.3
+Analysis of Scale-Aware State Fuser.
+In Table 5, we present the strategies for fusing the outputs of each TEMBA block, to learn the multi-scale temporal representations.
+In the fuser, we aim to first merge the representations and then have an interaction between the merged tokens to have a unified multi-scale representation.
+First, following [8], we concatenate the features and have a Temporal Convolution layer for interaction among the tokens (denoted by ‘Concat + Temporal Conv.’).
+It is observed that a linear interaction between the tokens after concatenation (denoted by ‘Concat + Linear Proj.’) outperforms the former as it does not impose any specific temporal constraints leading to an effective aggregation.
+Furthermore, we find that aggregation through summation of the features leads to better performance.
+Finally, a BiMamba interaction layer (‘Sum + BiMamba Int.’) after the aggregation leads to the optimal results, which demonstrates the ability of BiMamba to learn an effective state space that represents the scale-aware features from each Temba block.
+Finding 3: SSMs can fuse homogeneous features within their hidden state-space.
+Dilation in TCM and D-SSM.
+Here, we explore the effectiveness of dilation in the Temba block, particularly in TCM and D-SSM.
+Our baseline is a model which does not use dilation in either the temporal convolution in TCM or D-SSM.
+As illustrated in Table 7, we find that using dilation in either TCM or D-SSM  results in a substantial improvement over the baseline scenario, highlighting the importance of dilation to capture multi-scale token interactions without scaling the Temporal dimension (see Appendix).
+Furthermore, using dilation in both TCM and D-SSM  leads to an improvement of $21.4\%$ in Charades and $68.7\%$ increase in TSU.
+Finding 4: SSMs with dilation can capture long-term dependencies more effectively than their counterparts.
 4.3 Other Design Choices
 In this section, we present the design choices of the Temba blocks in MS-Temba, particularly the scanning strategy of the SSM and number of Temba blocks.
+Table 8: Effect of Number of Temba Blocks
+#Blocks TSU Charades
+1 38.6 30.8
+2 39.5 31.6
+3 42.0 32.3
+4 25.0 31.0
+Table 9: Impact of Scaling Temporal Dimension
+TMF TSU Charades
+Max Pooling 30.8 31.8
+Average Pooling 36.3 30.9
+Strided Conv 37.8 31.2
+No Scaling 42.0 32.3
 Scanning Strategy.
@@
 Further increasing the number of blocks leads the model to learn redundant representations.
-Table 6: Effect of Scanning Strategy in SSM
-Scanning TSU Ch
-Unidirectional 37.0 31.0
-Bidirectional 42.0 32.3
-Table 7: Effect of Number of Temba Blocks
-#Blocks TSU Charades
-1 38.6 30.8
-2 39.5 31.6
-3 42.0 32.3
-4 25.0 31.0
-(a) Comparison of Information Density of I3D and CLIP features of a sample of 15 frames averaged over 10 videos of Charades
-(b) Comparison of Information Density of I3D and CLIP features of a sample of 100 frames averaged over 10 videos of TSU
+Effect of Temporal Dimension Scaling.
+Table 9 examines the impact of scaling the temporal dimension across successive Temba Blocks in MS-Temba.
+Each block halves the temporal dimension relative to its predecessor.
+To maintain a consistent number of tokens in Scale-Aware State Fuser, we upsample the temporal dimension to match the number of input tokens.
+Temporal downsampling is explored using pooling methods (max pooling, average pooling) and strided convolutions in TCM.
+However, reducing the number of tokens leads to information loss in the Temba Blocks’ state space, degrading model performance.
+Finding 5: Multiscale temporal representations obtained through downsampling are suboptimal for Mamba, as SSMs perform more effectively when operating on the full temporal sequence of tokens.
+5 Conclusion
+In this work, we introduce the first Mamba-based approach for Temporal Action Detection (TAD).
+Our study shows that directly adapting Mamba from other tasks is ineffective for TAD.
+By strategically integrating convolutions for local feature extraction and introducing dilations in SSMs for long-range dependencies, we successfully tailor Mamba for TAD.
+Additionally, our method leverages SSMs to learn robust multi-scale representations, particularly benefiting from the high information density of foundation model features.
+MS-Temba achieves state-of-the-art performance on long-duration videos with only 10M parameters, significantly outperforming existing methods.
+These advantages position MS-Temba as a highly efficient and effective solution, ideal for edge deployment and real-world applications.
+Supplementary Material
+6 Overview
+The supplementary is categorized into the following parts:
+  • Section 7: Preliminaries of State Space Models
+  • Section 8: Comparison of CLIP and I3D Features
+  • Section 9: Training configurations for each Dataset
+  • Section 10: Qualitative Evaluation of MS-Temba
+  • Section 11: Limitations of LLMs in Temporal Action Detection
+  • Section 12: Effect of the number of Mamba Layers in MS-Temba
+  • Section 13: Architecture of MS-Temba
+7 Preliminaries of State Space Models
+State space models map continuous input sequences $x(t)\in\mathbb{R}$ to continuous outputs $y(t)\in\mathbb{R}$ using a hidden state $h(t)\in\mathbb{R}^{n}$.
+For the above mapping and updating of the hidden state, SSMs use a transition state matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ projection matrices $\mathbf{B}\in\mathbb{R}^{n\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times n}$.
+The update equations are given by
+$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t)$ (10)
+$y(t)$ $=\mathbf{C}h(t)$
+Previous works  [19, 18, 60], use discrete versions of the above system using a discretization parameter $\mathbf{\Delta}$ known as the step size.
+Following is an example of a discretization transformation using $\mathbf{\Delta}$, known as zero-order hold (ZOH):
+$\overline{\mathbf{A}}$ $=\exp(\mathbf{\Delta}\mathbf{A})$ (11)
+$\overline{\mathbf{B}}$ $=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp(\mathbf{\Delta}\mathbf{A})-\mathbf{I})\mathbf{\Delta}\mathbf{B}$
+This discretization of $\mathbf{A}$, $\mathbf{B}$ to $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ leads to the following updated state-space equations:
+$h_{t}$ $=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t}$ (12)
+$y_{t}$ $=\mathbf{C}h_{t}$
+Standard SSMs are typically designed to process 1D inputs in a linear sequence.
+However, [60, 27] introduce bidirectional scanning of tokens (illustrated in Figure 3), which are more effective for vision tasks.
+In this approach, the SSM operates on the input sequence $v$ and it’s reversed counterpart $v^{\prime}$, fusing the projected outputs for robust bidirectional sequence modeling.
+For brevity, we call it BiMamba for the remainder of the paper.
+This enhancement enables more effective representation learning, especially in domains requiring complex temporal dependencies.
+In this work, we extend the Mamba based architectures for understanding long untrimmed videos.
+8 Comparison of CLIP and I3D Features
 Figure 9: Comparison of Information Density of I3D and CLIP features in Charades and TSU
-6 Comparison of CLIP and I3D Features
 The superior performance of MS-Temba in action detection, leveraging features from foundation models instead of traditional 3D ConvNets, warrants the examination of the distinctions between features extracted by a CLIP backbone and and I3D backbone — widely adopted in prior state-of-the-art approaches[8, 7, 49].
@@
 Here, we use cosine similarity to measure the similarity between frame features.
-In Figure 10, we compare the information density of features extracted using a CLIP and an I3D backbone on a random sample of 15 frames for Charades (Fig.
-10) and 100 frames of TSU (Fig.
-10).
+In Figure 9, we compare the information density of features extracted using CLIP and I3D backbones.
+To ensure a unified representation across the data sets, the number of frames sampled - 15 for Charades and 100 for TSU - is normalized to the range [0,1].
 Figure 9 shows the information density averaged over 10 videos from the respective datasets.
@@
 We leave the exploration of this property of Mamba for other visual understanding tasks as future work.
-7 Conclusion
-In this work, we propose Multiscale Temporal Mamba (MS-Temba), leveraging the recent state-space model (SSM), Mamba, as a temporal encoder for action detection in untrimmed videos.
-It combines the strengths of SSMs and convolutions: convolutions model local information, while SSMs capture global dependencies.
-To enhance long-term relationship modeling, we introduce dilations in Mamba.
-This combination enables MS-Temba to learn temporal representations at multiple scales.
-We also utilize Mamba blocks to aggregate scale-specific information into a unified representation, which is then classified for action detection.
-MS-Temba achieves state-of-the-art performance on long-video datasets and matches the best-performing method on short-video datasets, using $88.5\%$ fewer parameters.
-Additionally, MS-Temba outperforms baseline models when leveraging visual features from foundation models.
-This work establishes a foundation for employing Mamba in efficient and effective long-video understanding.
-Acknowledgements
-We thank the members of the Genius Lab at UNC Charlotte for helpful discussions. This work is supported in part by the National Science Foundation (IIS-2245652).
-Supplementary Material
-8 Overview
-The supplementary is categorized into the following parts:
-  • Section 9: Training configurations for each Dataset
-  • Section 10: Effect of Temporal Dimension Scaling
-  • Section 11: Effect of the number of Mamba Layers in MS-Temba
-  • Section 12: Architecture of MS-Temba
+9 Training configurations for each Dataset
 Table 10: MS-Temba Training Settings
@@
 Total Epochs $140$ $100$ $100$
-9 Training configurations for each Dataset
 For all experiments, we pad the features of each input video to a fixed sequence length. Table 10 shows the configurations used for each dataset.
-Table 9: Impact of scaling Temporal Dimension in MS-Temba
-TMF TSU Charades
-Max Pooling 30.8 31.8
-Average Pooling 36.3 30.9
-Strided Conv 37.8 31.2
-No Scaling 42.0 32.3
-Figure 11: Impact of Number of Mamba Layers in TEMBA Block
+10 Qualitative Evaluation of MS-Temba
+Figure 10: Qualitative Evaluation of Action Detection Heatmap on an Example Video. refers to the case where MSTCT fails to detect any action class.
+Figure 10 presents a qualitative comparison of the predictions generated by MS-Temba and MS-TCT on a sample untrimmed video.
+Notably, MS-Temba demonstrates superior capability in detecting short-duration actions, such as Putting Something on a Table and Stirring Coffee/Tea, which MS-TCT fails to recognize, as well as long-duration actions like Walking, with greater accuracy.
+11 Limitations of LLMs in Temporal Action Detection
+Figure 11: Comparison of multimodal LLMs (MLLMs) and our approach on long untrimmed videos.
+While multimodal LLMs [29, 37] excel at capturing global spatio-temporal semantics for short clips, they often fail to identify the precise timestamps corresponding to frames of interest.
+TimeChat [41], despite being trained on timestamps of input videos, fails to detect concurrent actions in complex videos.
+Furthermore, TAD in long videos using MLLMs requires enormous computational resources (7B parameters), making them infeasible for edge devices.
+As shown in Figure 11, both VideoLLaVa [29] and TimeChat [41] struggle to identify dense concurrent actions in untrimmed videos.
 Table 11: MS-Temba architecture.
@@
 Fuser BiMamba $D_{s}$: 576 256$\times$576 256$\times$576
-10 Effect of Temporal Dimension Scaling
-In this section, we explore strategies for scaling the temporal dimensions in untrimmed videos to emulate the concept of hierarchical Mamba models.
-Similar temporal scaling technique is also employed in the state-of-the-art MS-TCT action detection model.
-Table 9 illustrates the effect of scaling the temporal dimension across successive Temba blocks in MS-Temba.
-The temporal dimension is halved in each block, compared to its predecessor.
-To maintain a consistent number of tokens in the TMF, we upsample the temporal dimension to match the number of input tokens.
-We explore this temporal scaling through downsampling the temporal dimension using pooling methods (maxpooling, average pooling) and strided convolutions in TLM.
-We observe that decreasing the number of tokens through temporal downsampling leads to loss of effective information in the state space of the Temba blocks thus leading to a degradation of the model’s performance.
 12 Effect of the number of BiMamba Layers in MS-Temba
+Figure 12: Impact of Number of Mamba Layers in TEMBA Block
 Previous state-of-the-art methods, relied on stacking multiple convolutional or Transformer layers in the Temporal Encoder to expand the receptive field and capture complex patterns at higher levels of abstractions.